need help with 2nd part of question.
Blood type AB is found in only 3% of the population†. If 285 people are chosen at random, find the probability of the following. (Use the normal approximation. Round your answers to four decimal places.)
(b) between 5 and 10 will have this blood type

Answer by Boreal(15235)

np=285*0.03=8.55, the mean
np(1-p)=variance=8.29
sqrt V=sd=2.88
want z between 4.5 and 10.5 using continuity correction factor
that would be low bound of (4.5-8.55)/2.88 or -4.1/2.88=-1.42 and upper of 1.95/2.88 or 0.68
the probability z is between those values is 0.6739